In an experiment to discover the acceleration due to gravity, a rock was dropped on the planet Zeus.
If it fell a distance of 15 m in 5 s, the acceleration of gravity on Zeus is
1
Expert's answer
2019-03-04T08:56:40-0500

Assuming that a rock was dropped without the initial velocity, one can use the following relation between the distance, acceleration and time of flight:


"H=\frac{g_{Z} t^2}{2},"

where gZ is the acceleration of gravity on Zeus. Deriving its value from the exression above, one can obtain:

"g_Z = \frac{2 H}{t^2}."

Substituting the numerical values, we get:


"g_Z = \frac{2 \cdot 15}{5^2} = 1.2 \, m/s^2"

Answer: 1.2 m/s2